# Spectrum and low-energy gap in triangular quantum spin liquid NaYbSe2

### v1

@@ Acknowledgments
The work by A.O.S., M.L., K.W., S.T., V.S.Z., C.D.B., J.E.M., and D.A.T. is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). The neutron scattering study used resources at the Spallation Neutron Source, a DOE Office of Science User Facility operated by the Oak Ridge National Laboratory. The work of P.L. and E.D. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. The work by H.Z. is supported by the U.S. Department of Energy under Grant No. DE-SC0020254. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. The work by Q.Z. and J. Ma was supported by the National Key Research and Development Program of China (Grant No. 2022YFA1402700). We acknowledge helpful discussions with Johannes Knolle.
@@ II Neutron Experiments
We measured the inelastic spectrum of NaYbSe2 using the $\sim 300$ mg co-aligned sample used in Ref. [30] mounted in a dilution refrigerator (no magnet was used in this experiment). We measured the $hhl$ inelastic scattering on the CNCS spectrometer [49] at Oak Ridge National Laboratory’s Spallation Neutron Source [50], measuring at $E_{i}=3.32$ meV, $1.55$ meV, and 1.0 meV, rotating 180 to map the neutron spectrum. We measured at $T=0.1$ K and 12 K for a background. The data are shown in main text Fig. 2, and were normalized to absolute units by normalizing the magnon mode measured in Ref. [30] to the nonlinear spin wave theory, such that the effective spin is 1/2.
@@
Figure 6: NaYbSe2 scattering with $E_{i}=1.0$ meV and $|\ell|\leq 1.0$ rlu. Panel (a) shows data along $hh$, panel (b) shows a constant $Q$ cut at $hh=(1/3,1/3)$. Intensity monotonically decreases with increasing energy, indicating a gapless spectrum to within ±$\pm 0.02$ meV.
@@
Figure 7: NaYbSe2 elastic scattering with $E_{i}=1.55$ meV at 0.1 K (a), 12 K (b), and 0.1 K -- 12 K (c), with an energy window ±$\pm 0.04$ meV. No static spin correlations are visible in the temperature-subtracted data, suggesting an absence of long-range magnetic order. Note that in the unsubtracted data, there are arcs of suppressed intensity from when the vertical plates of the sample holder are along the incident and scattered beams respectively, and absorption is much larger.
@@ III AC Calorimetry
Ac calorimetry measurements under hydrostatic pressure were performed in a piston-clamp pressure cell using Daphne oil 7373 as the pressure medium using the standard steady state technique [51]. The temperature oscillations were measured using an Au/0.07%Fe-chromel thermocouple, and a constantan meander was attached to the opposite side of the sample to apply heat. The heater power was varied between 25 nW and 5 $\rm\mu$W depending on the sample temperature. The measurement frequency was continuously adjusted to keep a constant phase relationship between the applied heat and the temperature oscillations on the thermocouple. For the lowest ($\leq$ 100 nW) powers and temperatures the frequency was fixed near 2 Hz because the signal was too small to continuously vary the frequency. Below 300 mK, it was not possible to find a frequency range where $f\Delta{}T_{ac}$ was constant. This indicates that the internal relaxation of the sample is likely slower than the relaxation rate to the bath. Nonetheless, the measurement would still be sensitive to phase transitions even in this temperature range.
In Fig. 11 the NaYbSe2 specific heat is compared to previously published KYbSe2 data [28]. The “bump” in $C/T$ is smaller in NaYbSe2 than in KYbSe2, while the specific heat below 300 mK is significantly larger in NaYbSe2. This indicates more of the density of states has shifted to low energies, which is consistent with the system being within a QSL phase. Note also, in main text Fig. 3 there is no significant missing entropy in NaYbSe2, which again is consistent with it being in a well-defined quantum ground state rather than a glassy frozen state.
@@ IV.1 Method
The ac susceptometer comprises a solenoidal coil to generate an ac magnetic field and a pair of sensing coils housed within it. The pair of sensing coils are wound in opposite directions, ensuring they possess equal mutual inductance in magnitude but opposite signs. Consequently, when two sensing coils are connected in series, the induced voltage across them becomes zero. The presence of a sample positioned in the center of one of the sensing coils induces a nonzero net voltage across the coils. This induced voltage is directly proportional to the change in magnetic flux passing through the sensing coil over time. More detailed information can be found in https://nationalmaglab.org/user-facilities/dc-field/measurement-techniques/ac-magnetic-susceptibility-dc/. This setup includes a nonzero background susceptibility. Based on our experience of running this setup for over ten years, we believe that the excessive susceptibility near zero magnetic field is due to coil background, although we did not perform a background measurement. The background in temperature scans is much smaller compared to the sample signal. Therefore, the susceptibility drop below 23 mK is due to the sample’s intrinsic behavior (confirmed by the absence of such a downturn in KYbSe2 data, see below). We used “Arbi. Unit” because of the background signal of the AC susceptometer.
@@ IV.2 Additional data
Figure 12 shows the temperature-dependent AC susceptibility at zero magnetic field up to higher temperatures than in the main text Fig. 4. Paramagnetic behavior is evident up to 500 mK, with no phase transitions visible.
@@
Figure 13 shows susceptibility to estimate the gap of the low-temperature drop in susceptibility, which we find to be 2.1 $\mu$eV.
Figure 13: AC susceptibility vs. temperature of NaYbSe2 with the AC field along the $a$-axis. The red curve represents the fitting with $\chi^{\prime}=Ae^{-\Delta/T}+\chi_{0}$. From this fit, we estimated the energy gap to be approximately 2.1 $\mu$eV.
@@
Finally, for comparison with NaYbSe2, Figure 15 shows the measured in-plane susceptibility of KYbSe2 (which was also measured in the same cryostat at the same time—and therefore the same temperature and field configurations—as the two NaYbSe2 crystals). Note the absence of a gap feature in the data, which follows a $1/T$ divergence to the lowest temperatures. Note also that the ordering transition is not visible in the data (which is admittedly somewhat noisy), again evidencing that the 23 mK downturn in NaYbSe2 is not from a magnetic ordering transition.
@@ V.3.1 NQS wave function
The NQS method utilizes an artificial neural network as a variational wave function to approximate the ground state of a target model [65]. In a system with $N$ spin-1/2 degrees of freedom, the Hilbert space can be spanned by the $S_{z}$ basis |$\ket{\sigma}=\ket{\sigma_{1},...,\sigma_{N}}$ with $\sigma_{i}=\,\uparrow$ or $\downarrow$. Similar to image recognition tasks in which the artificial neural network converts every image input to a probability, in quantum many-body problems the NQS converts every input basis |$\ket{\sigma}$ to a wave function amplitude $\psi_{\sigma}$. This gives the full quantum state as
|$\ket{\Psi}=\sum_{\sigma}\psi_{\sigma}\ket{\sigma}.$ (2)

### v2

@@ Acknowledgments
The work by A.O.S., M.L., K.W., S.T., V.S.Z., C.D.B., J.E.M., and D.A.T. is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). The neutron scattering study used resources at the Spallation Neutron Source, a DOE Office of Science User Facility operated by the Oak Ridge National Laboratory under IPTS-30453 and IPTS-28649. The work of P.L. and E.D. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. The work by H.Z. is supported by the U.S. Department of Energy under Grant No. DE-SC0020254. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. The work by Q.Z. and J. Ma was supported by the National Key Research and Development Program of China (Grant No. 2022YFA1402700). We acknowledge helpful discussions with Johannes Knolle.
@@ II Neutron Experiments
We measured the inelastic spectrum of NaYbSe2 using the $\sim 300$ mg co-aligned sample used in Ref. [30] mounted in a dilution refrigerator (no magnet was used in this experiment). We measured the hhlhhl inelastic scattering on the CNCS spectrometer [45] at Oak Ridge National Laboratory’s Spallation Neutron Source [46], measuring at $E_{i}=3.32$ meV, $1.55$ meV, and 1.0 meV, rotating 180 to map the neutron spectrum. We measured at $T=0.1$ K and 12 K for a background. The data are shown in main text Fig. 2, and were normalized to absolute units by normalizing the magnon mode measured in Ref. [30] to the nonlinear spin wave theory, such that the effective spin is 1/2 and is amenable to QFI calculation [47].
@@
Figure 6: NaYbSe2 scattering with $E_{i}=1.0$ meV and $|\ell|\leq 1.0$ rlu. Panel (a) shows data along $hh$, panel (b) shows a constant $Q$ cut at $hh=(1/3,1/3)$. Intensity monotonically decreases with increasing energy, indicating a gapless spectrum to within $\pm 0.02$ meV.
@@
Figure 7: NaYbSe2 elastic scattering with $E_{i}=1.55$ meV at 0.1 K (a), 12 K (b), and 0.1 K - 12 K (c), with an energy window $\pm 0.04$ meV. No static spin correlations are visible in the temperature-subtracted data, suggesting an absence of long-range magnetic order. Note that in the unsubtracted data, there are arcs of suppressed intensity from when the vertical plates of the sample holder are along the incident and scattered beams respectively, and absorption is much larger.
@@ III AC Calorimetry
[…]
In Fig. 11 the NaYbSe2 specific heat is compared to previously published KYbSe2 data [28]. The “bump” in $C/T$ is smaller in NaYbSe2 than in KYbSe2, while the specific heat below 300 mK is significantly larger in NaYbSe2. This indicates more of the density of states has shifted to low energies, which is consistent with the system being within a QSL phase. Note also, in main text Fig. 3 there is no significant missing entropy in NaYbSe2, which again is consistent with it being in a well-defined quantum ground state rather than a non-ergodic glassy frozen state.
@@ IV.1 Method
The ac susceptometer comprises a solenoidal coil to generate an ac magnetic field and a pair of sensing coils housed within it. The pair of sensing coils are wound in opposite directions, ensuring they possess equal mutual inductance in magnitude but opposite signs. Consequently, when two sensing coils are connected in series, the induced voltage across them becomes zero. The presence of a sample positioned in the center of one of the sensing coils induces a nonzero net voltage across the coils. This induced voltage is directly proportional to the change in magnetic flux passing through the sensing coil over time. More detailed information can be found in https://nationalmaglab.org/user-facilities/dc-field/measurement-techniques/ac-magnetic-susceptibility-dc/. This setup includes a nonzero background susceptibility. Based on our experience of running this setup for over ten years, we believe that the excessive susceptibility near zero magnetic field is due to coil background, although we did not perform a background measurement. The background in temperature scans is much smaller compared to the sample signal. Therefore, the susceptibility peak at 40 mK is due to the sample’s intrinsic behavior (confirmed by the absence of such a downturn in KYbSe2 data, see below). We used “Arbi. Unit” because of the background signal of the AC susceptometer.
To confirm the intrinsic nature of the 40 mK peak we also measured ultra-low temperature susceptibility down to 0.7 mK at the High B/T facility. This measurement used a much smaller excitation field of 0.01 Oe, but the peak is still present (albeit much less dramatic and at slightly higher temperature). Thus we associate the low temperature susceptibility downturn to intrinsic NaYbSe2 behavior. More detailed information can be found in https://nationalmaglab.org/user-facilities/high-b-t/magnets-instruments/8-tesla-bay-2/
@@ IV.2 Additional data
Figure 12 shows the temperature-dependent AC susceptibility at zero magnetic field up to higher temperatures than in the main text Fig. 4. Paramagnetic behavior is evident up to 500 mK, with no phase transitions visible. Note that with this higher excitation field, there appears to be sample heating which shifts the susceptibility downturn lower in temperature and makes it asymmetric (which we initially misinterpreted as a gap).
@@
Finally, for comparison with NaYbSe2, Figure 14 shows the measured in-plane susceptibility of KYbSe2 (which was also measured in the same cryostat at the same time—and therefore the same temperature and field configurations—as the two NaYbSe2 crystals). Note the absence of a downturn in the data, which follows a $1/T$ divergence to the lowest temperatures. Note also that the ordering transition is not visible in the data (which is admittedly somewhat noisy), consistent with the very subtle weak static magnetic order in this compound.
@@ V.3.1 NQS wave function
[…]